OpenStudy (anonymous):

The cube has a surface area of 24 cm2. What would the surface area of the prism be if each dimension were quadrupled?

OpenStudy (anonymous):

576?

OpenStudy (anonymous):

Hi Welcome to Openstudy! \(\sf Surface\ Area= 6a^2\) If we quadruple the dimension it will be like: \(\sf Surface\ Area"= 6(4a)^2\\Surface\ Area"= 6(16a^2)\\Suraface\ area"= 96a^2\) "a" is the dimension which is 2 cm I got a different answer. Can you show me how did you get 576?

OpenStudy (anonymous):

whats your answer

OpenStudy (anonymous):

i did 24*24

OpenStudy (anonymous):

a= 2 cm \(\sf Surface\ Area= 6a^2\) -> if you plug 2 cm here you'll get 24cm^2 If we quadruple the dimension it will be like: \(\sf Surface\ Area"= 6(4a)^2\) <- if you evaluate this, you'll get the right answer.

OpenStudy (anonymous):

2^4 = ? @i-love-skittles-lol-1

OpenStudy (anonymous):

what is 2^4 ? its 16

OpenStudy (anonymous):

, and the formula for the surface area of a cube is: SA = 6a^2 Where 'a' is the side length, plug it in.

OpenStudy (anonymous):

6(16)2 are you getting it?

OpenStudy (anonymous):

@Hannah_Waller May I ask why would you do 2^4 ? The question says "each dimension were quadrupled" so if we solve the surface area without the formula it goes like this: Dimension will be: 2 x 4 = 8 so each dimension is 8 A cube has 6 equal sides and each side of a cube is a square. Formula for the area of square is \(\sf a^2\) so the S.A. would be: \(\sf a^2 + a^2+ a^2+ a^2+a^2+a^2\) where the value of a this time is 8 cm not 16 cm.